# Evidence of Scaling Regimes in the Hopfield Dynamics of Whole Brain Model

### v2

@@
Evidence of Scaling Regimes in the Hopfield Dynamics of Whole Brain Model
Giorgio Gosti,1,2∗ Sauro Succi,1,3 Giancarlo Ruocco1
1Center for Life Nano- and Neuro-Science, Istituto Italiano di Tecnologia,
Viale Regina Elena 201, I-00161, Rome, Italy
2Istituto di Scienze del Patrimonio Culturale, Consiglio Nazionale delle Ricerche,
00010, Montelibretti (RM), Italy
3Istituto per le Applicazioni del Calcolo del Consiglio Nazionale delle Ricerche,
via dei Taurini 19, I-00185, Rome, Italy
4Dipartimento di Fisica, Università di Roma “La Sapienza”,
P.le Aldo Moro 5, I-00185, Roma, Italy
To whom correspondence should be addressed; E-mail: giorgio.gosti@cnr.it
@@
It is shown that a Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently introduced by Deco et al., according to which the process of information transfer within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows, although with a more singular exponent, $1/2$ instead of $2/3$. Both models employ a coupling strength which decays exponentially with the Euclidean distance between the nodes, but their mathematical nature is very different, Hopf oscillators versus a Hopfield neural network, respectively. Hence, their convergence for the same data parameters, suggests an intriguing robustness of the scaling picture. The present analysis shows that the Hopfield model brain remains functional by removing links above about five decay lengths, corresponding to about one sixth of the size of the global brain. This suggests that, in terms of connectivity decay length, the Hopfield brain functions in a sort of intermediate “turbulent liquid”-like state, whose essential connections are the intermediate ones between the connectivity decay length and the global brain size. Finally, the scaling exponents are shown to be highly sensitive to the value of the decay length, as well as to the number of brain parcels employed. As a result, any quantitative assessment regarding the specific nature of the scaling regime must be taken with great caution.
Introduction
@@
In a recent paper  (?), it was argued that the information transfer within the human brain may proceed in close analogy with the mechanisms which govern mass and energy transport in turbulent fluids. The analogy refers to the emergence of correlated spacetime patterns which would facilitate and possibly optimize information transfer across the brain, in the “same” way as fluid turbulence enhances energy transport across fluids. This is interesting in many respects, not least the prospects for future neurocomputers  (?).
@@
Based on an extended dynamical system of oscillators (Hopf model), supplemented by large-scale neuro-imaging empirical data on the connections between different brain areas, the authors of  (?) numerically showed that the statistical correlations of human brain signals may exhibit a scaling regime similar to the one observed in turbulent fluids.
@@
At a closer scrutiny, though, the analogy does not appear to be quantitative, meaning that the spatial scaling exponents of the second order structure function $S_{2}(d)$=$\langle(O(d)-O(0))^{2}\rangle$$\sim d^{\alpha}$, $O$ being a suitable order parameter, are not the same, about $\alpha$$\sim$1/2 for the Hopf model of the human brain against $\alpha$=2/3 for fluid turbulence. Hence “brain turbulence” appears to be more singular than fluid turbulence.
@@
The recurrent Hopfield neural network model is a well-established model for small-scale neuronal networks  (?, ?, ?). It is a neural network model based on binary McCulloch-Pitts neurons  (?, ?, ?, ?) in which every processing unit ($i$) is connected to all the other ones ($j$) through a set of weights ($J_{ij}$). Most of the literature considers fully connected and symmetric Hopfield networks since this simplifies the mathematical analysis of the model, although recently the investigation of diluted networks  (?, ?, ?, ?, ?) and asymmetric networks  (?, ?, ?, ?, ?, ?, ?, ?) was introduced.
@@
Usually, a recurrent Hopfield neural network is trained to store patterns or memories as steady-state attractors with the Hebbian prescription. In this case, the coupling matrix $J_{ij}$ takes the form of the sum of diadiics, and therefore it is symmetric and fully connected. The maximum storage capacity of such a network is found to scale linearly with the number of nodes $N$. The capacity can be increased by the process of learning (Hebbian learning) whereby the simultaneous activation of two nodes reinforces their connection. This is inspired by biological neural networks where the simultaneous activation of neurons leads to increments in synaptic strength  (?), and it allows the network to store a larger number of patterns  (?). Yet, this quantity ($C$) still remains linear with $N$: $C\sim N$, being the proportionally factor of the order of unity. Lately, researchers have also investigated how unlearning methods can improve the stability of the stored patterns  (?, ?, ?).
@@
Interestingly, it has been recently demonstrated both numerically  (?, ?) and theoretically  (?, ?) that random, non-Hebbian, coupling matrices have a storage capacity much larger than the Hebbian ones. Indeed the storage capacity $C$ of random matrices scales exponentially with $N$: $C\sim e^{\Sigma N}$ being $\Sigma$ the complexity. The latter quantity is found to be maximized for asymmetric and diluted matrices, with values that match the connectivity found in the hippocampus areas.
@@
The approach outlined in reference  (?) diverges from both the Hebbian and random connectivity models. It relies on a coupling matrix, which is derived from experimental observations of the connections between human brain regions. This approach assumes that the coupling strengths are all positive, deterministic, and specifically proportional to a given function of the distances between nodes. $J_{ij}=e^{-d_{ij}/\delta}$, $d_{ij}$ being the Euclidean distance between nodes $i$ and $j$ in the network, and $\delta$ a decay length, which, in  (?) are kept fix to match the empirical HCP dMRI tractography of the human brain: $\delta=5.55$ $\mathrm{m}\mathrm{m}$.
@@
At variance with the latter approach, where the dynamics is based on the Hopf oscillator model, here we use the same coupling matrix of  (?), but employ the Hopfield dynamics.
@@
The use of Hopfield networks to investigate the collective dynamics of the brain activity is not new. Indeed,  (?), and  (?) used Hopfield models (or Ising-Spin models with Glauber dynamics) to study the complex landscape of brain activation dynamics. These schemes do not interpret the nodes of the Hopfield network as single spiking neurons but rather as aggregates of neural mass (parcel), namely mesoscale brain regions switching between active and rest states. In the active states many neurons simultaneously fire while in the rest states most neurons are silent. These models do not aim at establishing a rigorous connection between the microscopic description of the brain and a corresponding mesoscopic one, but posit instead a physically-informed heuristic picture of the collective brain dynamics.
Broadly speaking, the Hopf model is based on the idea of reproducing the empirically observed information transfer between areas that takes place through (partial) synchronization of collective oscillations. In this framework, the natural collective variables describing the global brain dynamics are mesoscale oscillators obeying the Hopf dynamics.
@@
The Hopfield model used in this work draws inspiration from a very different idea, namely to carry over the discontinuous divide between active and rest states from the microscale level of single neurons to the mesoscale level of neural parcels, as discussed above. Formally, this amounts to assuming that coarse-graining of the Hopfield model would leave the model invariant but with a different set of weights, in our case an exponential decay. Such model proves useful to show that the dynamical repertoire of brain activity composed of sudden avalanches and bursts, can be described in terms of critical systems perturbed by noise  (?, ?).
@@
Due to the significant difference between the fundamental assumptions behind these two models, inspecting their commonolaties, as well as their points of departure, starting from the same empirical connectivity information, is of decided interest to improve the mesoscale description of the global brain dynamics,
@@
ii) At the $\delta$ value investigated in  (?) ($\sim 5.5$ mm) and for $N=1000$, we found $\alpha\sim 2/5$, close to the one reported by Deco et al.  (?) [^1] In Ref.  (?) the used parameter is the inverse of the decay scale, $\lambda=1/\delta$.;
iii) For $N=1000$ the scaling exponent $\alpha$ shows a sigmoid-like trend with the decay length $\delta$. Importantly, for $N=1000$, the functional relation $\alpha(\delta)$ shows a steep increase in the region around $\delta\sim 5.55$ mm, the value indicated in  (?) as the physiologically relevant one from the analysis of empirical HCP dMRI tractography with a $N=1000$ parcellation. This means that the specific value of the decay length has a significant impact on the scaling exponent; for instance, a slight increase to $\delta$=5.88 mm would yield a scaling exponent slightly above $2/3$, the value associated with homogenoeus incompressible turbulence. Conversely, a slight decrease to $\delta$=5 mm, would yield $\alpha$$\sim$0.1, a much less organized regime.
@@
Discrete-time Hopfield Recurrent Neural Networks
Following  (?, ?, ?, ?, ?, ?, ?, ?), we consider a network of $N$-binary nodes that can be either firing or silent.The state of each node is a binary variable $s_{i}$, with $i=1,2,\ldots,N$, such that $s_{i}=-1$ or $s_{i}=1$, so that the network state is represented by a binary string $\mathbf{s}=(s_{0},s_{1},\ldots,s_{N})$. Neurons interact through the connectivity matrix $\mathbf{J}$, with matrix elements $J_{ij}$ following the same exponential dependence $J_{ij}=e^{-d_{ij}/\delta}$ used in  (?). We assume a discrete time $t$, and, at each time step $t$, the evolution of the neuron state $s_{i}(t)$ is given by the following non-linear dynamic equation,
@@
We determined the couplings matrix $J_{ij}$ using Schaefer’s cerebral cortical parcellation atlas as in  (?). This parcellation was obtained from the analysis of a large dataset of neural activity dynamics with a sample size of 1489  (?). This dataset is publicly available, we used Schaefer’s parcellations with $N$ ranging from $200$ to $1000$. The Schaefer’s parcel centroids are mapped onto the MNI152 volumetric space [^2] Github repository dataset url  (?): https://github.com/ThomasYeoLab/CBIG/tree/master/stable\_projects/brain\_parcellation/Schaefer2018\_LocalGlobal/Parcellations/MNI/Centroid\_coordinates. The code for reproducing the simulations can be found in the git hub repository https://github.com/ggosti/HopBrain.
The Structure Factor $S_{2}(d)$
We used  (?) adaptation of Kolmogorov’s concept of structure functions. In turbulence, the Kolmogorov’s structure factor applies to the longitudinal velocity but in our case it represents the activity of the parcel as measured through the BOLD signal in a fMRI experiment.
@@
where the function ${\cal{N}}(d)$ represents the number of $i,j$ pairs at distance $d$. The notation $d_{ij}\in d$ indicates that the distance between the nodes must be “close enough” to $d$. To this aim, it should be kept in mind that in Schaefer’s cerebral cortical parcellation atlas, $d$ was recorded with 2 mm precision, thus there are multiple $i,j$ pairs with the same $d$. Therefore we can bin the existing nodes’ distance and map them to the natural numbers, so as to use of the Kronecker delta $\Delta(ij|d)$, which takes value $1$ if $d_{ij}$ is in the bin identified by $d$, and is 00 otherwise.
@@
The function $B(d)$ relates to the structure factor as follows  (?):
@@
The Hopfield dynamics is deterministic, and since the phase space of the nodes’ states is finite, the asymptotic trajectories must necessarily be periodic. Therefore, as also discussed in  (?), Hopfield Recurrent Neural Networks, given an initial state, consistently end up into a specific steady state (“attractor”). The steady states are deterministic attractors composed of either limit cycles formed by repeating sequence of states, or a single stationary state. Given that in our case the couplings $J_{ij}$ are symmetric, we expected theoretically, and experimentally confirmed that the steady states are always composed of a single fixed point. In the absence of a further external input (“external stimuli”), and given sufficient time, the network gets stuck in an attractor and the whole activity is constant in time, hence no time summation was taken in the previous equations. This is of course a very crude representation of the brain dynamics, which receives continuously “internal” and “external” inputs, that perturb the system to constantly explore a landscape shaped by the collection of steady states. Indeed,  (?), and  (?) show how the whole set of attractors allows us to reproduce the dynamical repertoire and the functional correlations of the brain activation dynamics.
Given different values of the decay length $\delta$, we have run the Hopfied model for a system composed by $N=1000$ parcels (with pairwise distances consistent with the Schaefer’s cerebral cortical parcellation atlas  (?)). The structure functions have then been calculated by following the system until the fixed points were reached, and by averaging over $N_{r}$=1000 randomly chosen initial states. As examples, figure 1 shows the structure factor $S_{2}(d)$ for respectively $\delta$=5.55 mm (Fig. 1A) and $\delta$=5.88 mm (Fig. 2B). The black points with the error bars represent the average $S(d)$ computed over 1000 different realizations. The dashed green line corresponds to $\alpha$=2/3, the orange line corresponds to $\alpha$=1/2, and the blue line corresponds to $\alpha$=2/5. The $\alpha$=1/2 value (orange line) is the value estimated in  (?). The $\alpha$=2/3 is the exponent associated with a turbulence. Due to the significant level of statistical fluctuations, we have estimated the average exponent in two different ways, first as the correct ensemble average $\langle S_{2}(d)\rangle$$\sim d^{\alpha}$ and also as an arithmetic average of the single realization exponents, $\overline{\alpha}$=$N_{r}^{-1}\sum_{r=1}^{N_{r}}\alpha_{r}$. Even though the latter is not correct, the fact that both yield very similar values attests to the robustness of the average exponent, notwithstanding sizeable fluctuations. For consistency with  (?), we estimate the scaling exponent $\alpha$ fitting the points in figure 1 in the interval 2.7 - 33.1 mm and we obtained $\alpha\approx$2/5 for $\delta$=5.55 mm, and $\alpha\approx$2/3 for $\delta$=5.88 mm.
@@
We repeated the previously depicted procedure for a sequence of values of the decay length in the range $[3-10]$ mm and measured the associated scaling exponents. The values of $\alpha(d)$ for different realizations are shown in Fig. 2 as blue points. The code stand instructions to reproduce the simulation and the 3D model files (gltf) are available at the repository https://github.com/ggosti/HopBrain. This code also shows how to generate 3D maps for the Schaefer parcellation’s nodes activation states inspection, and how to export gltf files that can be viewed in immersive environments using open source webXR apps such as Aton which are part of the H2IOSC open cloud  (?, ?) (Fig. 2B-D).

### v3

@@
Abstract
@@
It is shown that a Hopfield recurrent neural network exhibits a scaling regime, whose specific exponents depend on the number of parcels used and the decay length of the coupling strength. This scaling regime recovers the picture introduced by Deco et al., according to which the process of information transfer within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows, although with a more singular exponent, $1/2$ instead of $2/3$. Both models employ a coupling strength which decays exponentially with the Euclidean distance between the nodes, informed by experimentally derived brain topology. Nevertheless, their mathematical nature is very different, Hopf oscillators versus a Hopfield neural network, respectively. Hence, their convergence for the same data parameters, suggests an intriguing robustness of the scaling picture. Furthermore, the present analysis shows that the Hopfield model brain remains functional by removing links above about five decay lengths, corresponding to about one sixth of the size of the global brain. This suggests that, in terms of connectivity decay length, the Hopfield brain functions in a sort of intermediate “turbulent liquid”-like state, whose essential connections are the intermediate ones between the connectivity decay length and the global brain size. The evident sensitivity of the scaling exponent to the value of the decay length, as well as to the number of brain parcels employed, leads us to take with great caution any quantitative assessment regarding the specific nature of the scaling regime.
keywords:
Hopfield Network , Whole-Brain Model , Neural Mass Model , Scaling Regimes
journal: Neural Networks
[…]
[a]organization=Center for Life Nano- and Neuro-Science, Istituto Italiano di Tecnologia,addressline=Viale Regina Elena 201, city=Rome, postcode=I-00161, country=Italy
[…]
[b]organization=Istituto di Scienze del Patrimonio Culturale, Consiglio Nazionale delle Ricerche,addressline=Strada della Neve s.n.c., Via Salaria km 29.300, city=Montelibretti (RM), postcode=I-00010, country=Italy
[…]
[c]organization=Istituto per le Applicazioni del Calcolo del Consiglio Nazionale delle Ricerche,addressline=via dei Taurini 19, city=Roma, postcode=I-00185, country=Italy
[…]
[d]organization=Dipartimento di Fisica, Universitá di Roma “La Sapienza”,addressline=P.le Aldo Moro 5, city=Roma, postcode=I-00185, country=Italy
{highlights}
The mesoscale Hopfield model of the global brain collective dynamics presents a scaling regime.
It recovers the scaling picture recently introduced by Deco et al. 2020 based on a Hopf oscillator model. The near-quantitative convergence of the two models points to an intriguing robustness of the scaling picture.
A scaling regime implies that the brain functions in a sort of intermediate “turbulent liquid”-like state.
It also shows that the scaling exponent associated with such collective patterns is highly sensitive to the decaying length of the brain connectivity, as well as to the number of brain parcels.
The global brain Hopfield model is shown to remain functional upon removing connections above five decay length scales.
1 Introduction
@@
In a recent paper, Deco and Kringelbach (2020) argued that the information transfer within the human brain may proceed in close analogy with the mechanisms which govern mass and energy transport in turbulent fluids. The analogy refers to the emergence of correlated spacetime patterns which would facilitate and possibly optimize information transfer across the brain, in the “same” way as fluid turbulence enhances energy transport across fluids. This is interesting in many respects, not least the prospects for future neurocomputers (Schuman et al., 2022).
@@
Based on an extended dynamical system of oscillators (Hopf model), supplemented by large-scale neuro-imaging empirical data on the connections between different brain areas, Deco and Kringelbach (2020) numerically showed that the statistical correlations of human brain signals may exhibit a scaling regime similar to the one observed in turbulent fluids.
@@
At a closer scrutiny, though, the analogy does not appear to be quantitative, meaning that the spatial scaling exponents of the second order structure function $S_{2}(d)$=$\langle(O(d)-O(0))^{2}\rangle$$\sim d^{\alpha}$, $O$ being a suitable order parameter, are not the same, about $\alpha$$\sim$1/2 for the Hopf model of the human brain against $\alpha$=2/3 for fluid turbulence (Benzi et al., 1993). Hence “brain turbulence” appears to be more singular than fluid turbulence.
@@
The recurrent Hopfield neural network model is a well-established model for small-scale neuronal networks (Amari, 1972; Little, 1974; Hopfield, 1982) and a particularly minimal discrete-time recurrent neural network model (Grossberg, 1967; Rumelhart and McClelland, 1986). It is a neural network model based on binary McCulloch-Pitts neurons (Amit et al., 1985; Brunel, 2016; Hillar and Tran, 2018; Hillar et al., 2021) in which every processing unit ($i$) is connected to all the other ones ($j$) through a set of weights ($J_{ij}$). Most of the literature considers fully connected and symmetric Hopfield networks since this simplifies the mathematical analysis of the model, although recently the investigation of diluted networks (Stauffer et al., 2003; Brunel, 2016; Kim et al., 2017; Folli et al., 2018; Leonetti et al., 2020) and asymmetric networks (Gopalsamy and zhong He, 1994; Xu et al., 1996; Chen and Amari, 2001; Franca and Yang, 2000; Zheng et al., 2010; Szedlak et al., 2014; Folli et al., 2018; Leonetti et al., 2020) was introduced.
@@
Usually, a recurrent Hopfield neural network is trained to store patterns or memories as steady-state attractors with the Hebbian prescription. In this case, the coupling matrix $J_{ij}$ takes the form of the sum of dyadics, and therefore it is symmetric and fully connected. The maximum storage capacity of such a network is found to scale linearly with the number of nodes $N$. The capacity can be increased by the process of learning (Hebbian learning) whereby the simultaneous activation of two nodes reinforces their connection. This is inspired by biological neural networks where the simultaneous activation of neurons leads to increments in synaptic strength (Hebb, 1949), and it allows the network to store a larger number of patterns as steady-states (Amit et al., 1985). Yet, this quantity ($C$) remains linear with $N$: $C\sim N$, being the proportionality factor of the order of unity. Lately, researchers have also investigated how unlearning methods can improve the stability of the stored patterns Fachechi et al. (2019); Benedetti et al. (2022); Benedetti and Ventura (2023); Agliari et al. (2019, 2024). It is worth emphasizing that a vast literature, tracing back to the seminal papers by Gardner (1986) and Hopfield et al. (1983), treat the stationary points of the Hopfield dynamics as “memories”, but in the context of whole brain dynamics, and in our work, we will not consider this interpretation.
@@
Interestingly, it has been recently demonstrated both numerically (Folli et al., 2018; Gosti et al., 2019) and theoretically (Hwang et al., 2019, 2020) that random, non-Hebbian, coupling matrices have a number of steady-states much larger than the Hebbian ones. Indeed the number of steady-states $C$ of random matrices scales exponentially with $N$: $C\sim e^{\Sigma N}$ being $\Sigma$ the complexity. The latter quantity is found to be maximized for asymmetric and diluted matrices, with values that match the connectivity found in the hippocampus areas.
@@
The approach outlined by Deco and Kringelbach (2020) diverges from both the Hebbian and random connectivity models. It relies on a coupling matrix, which is derived from experimental observations of the connections between human brain regions. This approach assumes that the coupling strengths are all positive, deterministic, and specifically proportional to a given function of the distances between nodes. $J_{ij}=e^{-d_{ij}/\delta}$, $d_{ij}$ being the Euclidean distance between nodes $i$ and $j$ in the network, and $\delta$ a decay length, which, in Deco and Kringelbach (2020) are kept fix to match the empirical HCP dMRI tractography of the human brain: $\delta=5.55$ $\mathrm{m}\mathrm{m}$.
@@
At variance with the latter approach, where the dynamics is based on the Hopf oscillator model, here we use the same coupling matrix of Deco and Kringelbach (2020), but employ the Hopfield dynamics.
@@
The use of Hopfield networks to investigate the collective dynamics of the brain activity is not new (Trappenberg, 2022). Indeed, Deco et al. (2012), and Golos et al. (2015) used Hopfield models (or Ising-Spin models with Glauber dynamics) as a neural mass model to study the complex landscape of brain activation dynamics. Neural mass models schemes do not interpret the nodes of the Hopfield network as single spiking neurons but rather as aggregates of neural mass (parcel), namely mesoscale brain regions switching between active and rest states. In the active states, many neurons simultaneously fire while in the rest states most neurons are silent. These models do not aim at establishing a rigorous connection between the microscopic description of the brain and a corresponding mesoscopic one, but posit instead a physically-informed heuristic picture of the collective brain dynamics.
@@
Broadly speaking, the Hopf model is based on the idea of reproducing the empirically observed information transfer between areas that takes place through (partial) synchronization of collective oscillations. In this framework, the natural collective variables describing the global brain dynamics are mesoscale oscillators obeying the Hopf dynamics. The Hopfield model used in this work draws inspiration from a very different idea, namely to carry over the discontinuous divide between active and rest states from the microscale level of single neurons to the mesoscale level of neural parcels, as discussed above. Formally, this amounts to assuming that coarse-graining of the Hopfield model would leave the model invariant but with a different set of weights, in our case an exponential decay. Such model proves useful to show that the dynamical repertoire of brain activity composed of sudden avalanches and bursts, can be described in terms of critical systems perturbed by noise (Deco et al., 2012; Golos et al., 2015).
@@
Due to the significant difference between the fundamental assumptions behind these two models, inspecting their commonalities, as well as their points of departure, starting from the same empirical connectivity information, is of decided interest to improve the mesoscale description of the global brain dynamics,
@@
ii) At the $\delta$ value investigated in Deco and Kringelbach (2020) ($\sim 5.5$ mm) and for $N=1000$, we found $\alpha\sim 2/5$, close to the one reported by Deco and Kringelbach (2020) [^1] In Ref. Deco and Kringelbach (2020) the used parameter is the inverse of the decay scale, $\lambda=1/\delta$.;
iii) For $N=1000$ the scaling exponent $\alpha$ shows a sigmoid-like trend with the decay length $\delta$. Importantly, for $N=1000$, the functional relation $\alpha(\delta)$ shows a steep increase in the region around $\delta\sim 5.55$ mm, the value indicated in Deco and Kringelbach (2020) as the physiologically relevant one from the analysis of empirical HCP dMRI tractography with a $N=1000$ parcellation. This means that the specific value of the decay length has a significant impact on the scaling exponent; for instance, a slight increase to $\delta$=5.88 mm would yield a scaling exponent slightly above $2/3$, the value associated with homogenoeus incompressible turbulence. Conversely, a slight decrease to $\delta$=5 mm, would yield $\alpha$$\sim$0.1, a much less organized regime.
@@
2 Discrete-time Hopfield Recurrent Neural Networks
Following (Amari, 1972; Little, 1974; Hopfield, 1982; Amit et al., 1985; Folli et al., 2017; Gosti et al., 2019; Leonetti et al., 2020; Gosti et al., 2024), we consider a network of $N$-binary nodes that can be either firing or silent. The state of each node is a binary variable $s_{i}$, with $i=1,2,\ldots,N$, such that $s_{i}=-1$ or $s_{i}=1$, so that the network state is represented by a binary string $\mathbf{s}=(s_{0},s_{1},\ldots,s_{N})$. Neurons interact through the connectivity matrix $\mathbf{J}$, with matrix elements $J_{ij}$ following the same exponential dependence $J_{ij}=e^{-d_{ij}/\delta}$ used in (Deco and Kringelbach, 2020). We assume a discrete time $t$, and, at each time step $t$, the evolution of the neuron state $s_{i}(t)$ is given by the following non-linear dynamic equation,
@@ 2 Discrete-time Hopfield Recurrent Neural Networks
We determined the couplings matrix $J_{ij}$ using Schaefer’s cerebral cortical parcellation atlas as in (Deco and Kringelbach, 2020). This parcellation was obtained from the analysis of a large dataset of neural activity dynamics with a sample size of 1489 (Schaefer et al., 2018). This dataset is publicly available. We used Schaefer’s parcellations with $N$ ranging from $200$ to $1000$. The Schaefer’s parcel centroids are mapped onto the MNI152 volumetric space [^2] Github repository dataset url Schaefer et al. (2018): https://github.com/ThomasYeoLab/CBIG/tree/master/stable_projects/brain_parcellation/Schaefer2018_LocalGlobal/Parcellations/MNI/Centroid_coordinates. The code for reproducing the simulations can be found in the git hub repository https://github.com/ggosti/HopBrain.
3 The Structure Factor $S_{2}(d)$
We used Deco and Kringelbach (2020) adaptation of Kolmogorov’s concept of structure functions. In turbulence, the Kolmogorov’s structure factor applies to the longitudinal velocity but in our case it represents the activity of the parcel as measured through the BOLD signal in a fMRI experiment.
@@ 3 The Structure Factor $S_{2}(d)$
where the function ${\cal{N}}(d)$ represents the number of $i,j$ pairs at distance $d$. The notation $d_{ij}\in d$ indicates that the distance between the nodes must be “close enough” to $d$. To this aim, it should be kept in mind that in Schaefer’s cerebral cortical parcellation atlas (Schaefer et al., 2018), $d$ was recorded with 2 mm precision, thus there are multiple $i,j$ pairs with the same $d$. Therefore we can bin the existing nodes’ distance and map them to the natural numbers, so as to use of the Kronecker delta $\Delta(ij|d)$, which takes value $1$ if $d_{ij}$ is in the bin identified by $d$, and is 00 otherwise.
@@
The function $B(d)$ relates to the structure factor as follows (Deco and Kringelbach, 2020):
@@
The Hopfield dynamics is deterministic, and since the phase space of the nodes’ states is finite, the asymptotic trajectories must necessarily be periodic. Therefore, as also discussed in (Folli et al., 2017), Hopfield Recurrent Neural Networks, given an initial state, consistently end up into a specific steady state (“attractor”). The steady states are deterministic attractors composed of either limit cycles formed by repeating sequence of states, or a single stationary state. Given that in our case the couplings $J_{ij}$ are symmetric, we expected theoretically, and experimentally confirmed that the steady states are always composed of a single fixed point. In the absence of a further external input (“external stimuli”), and given sufficient time, the network gets stuck in an attractor and the whole activity is constant in time, hence no time summation was taken in the previous equations. This is of course a very crude representation of the brain dynamics, which receives continuously “internal” and “external” inputs, that perturb the system to constantly explore a landscape shaped by the collection of steady states. Given this context, Deco et al. (2012), and Golos et al. (2015) show how the whole set of attractors allows us to reproduce the dynamical repertoire and the functional correlations of the brain activation dynamics.
@@
Figure 1: Structure factor $S_{2}(d)$ for $N=1000$ nodes and $N_{r}=1000$ realizations respectively at $\delta=5.55$ mm (A), and $\delta=5.88$ mm (B). The values $S_{2}(d)$ were binned over equal sized intervals. The dashed lines indicate different $\alpha$ values: $\alpha=2/3$ which corresponds to turbulence, $\alpha=1/2$, the value obtained in (Deco and Kringelbach, 2020), and $\alpha=2/5$. We estimate the scaling exponent $\alpha$ with a linear regression of $\log(S(d))\sim\alpha\log(d)+\beta$ in the interval range $[2.7-33.1]$ mm, as used in (Deco and Kringelbach, 2020), and we obtain $\alpha\approx 2/5$ for $\delta=5.55$ mm, and $\alpha\approx 2/3$ for $\delta=5.88$ mm.
Given different values of the decay length $\delta$, we have run the Hopfied model for a system composed by $N=1000$ parcels (with pairwise distances consistent with the Schaefer’s cerebral cortical parcellation atlas (Schaefer et al., 2018)). The structure functions have then been calculated by following the system until the fixed points were reached, and by averaging over $N_{r}$=1000 randomly chosen initial states. As examples, figure 1 shows the structure factor $S_{2}(d)$ for respectively $\delta$=5.55 mm (Fig. 1A) and $\delta$=5.88 mm (Fig. 1B). The black points with the error bars represent the average $S(d)$ computed over 1000 different realizations. The dashed green line corresponds to $\alpha$=2/3, the orange line corresponds to $\alpha$=1/2, and the blue line corresponds to $\alpha$=2/5. The $\alpha$=1/2 value (orange line) is the value estimated in (Deco and Kringelbach, 2020). The $\alpha$=2/3 is the exponent associated with a turbulence. Due to the significant level of statistical fluctuations, we have estimated the average exponent in two different ways, first as the correct ensemble average $\langle S_{2}(d)\rangle$$\sim d^{\alpha}$ and also as an arithmetic average of the single realization exponents, $\overline{\alpha}$=$N_{r}^{-1}\sum_{r=1}^{N_{r}}\alpha_{r}$. Even though the latter is not correct, the fact that both yield very similar values attests to the robustness of the average exponent, notwithstanding sizeable fluctuations. For consistency with (Deco and Kringelbach, 2020), we estimate the scaling exponent $\alpha$ fitting the points in figure 1 in the interval 2.7 - 33.1 mm and we obtained $\alpha\approx$2/5 for $\delta$=5.55 mm, and $\alpha\approx$2/3 for $\delta$=5.88 mm.
Figure 2: A) The blue line shows the average slope of $\alpha=\log S(d)/\log d$ for different $\delta$s, given the connectivity matrix $J_{ij}=\exp{(-d_{i,j}/\delta)}$. The horizontal dotted lines correspond to $\alpha_{RW}=1$ and $\alpha_{T}=2/3$, corresponding to random walk and homogeneous incompressible turbulence, respectively, while the dot-dashed horizontal line corresponds to Deco and Kringelbach (2020) “brain” value $\alpha_{D}=1/2$. The orange line shows the function $\alpha(\delta)$ with a connectivity matrix obtained by randomly shuffling all the weights $J_{ij}$, which results into a basically uncorrelated signal, indicating the importance of the metric structure of the weights. The plot stops at $\delta=10$ mm, a region where the signal is highly correlated ($\alpha\sim 1.2$), but still below the smooth regime marked by $\alpha=2$. Conversely, a brain below $4$ mm, would exhibit “chaotic” behaviour, with scaling exponents close to zero, i.e. no scaling regime. The bottom panels show the neuron activity mapped on the Schaefer’s cerebral cortical parcellation atlas (Schaefer et al., 2018) coordinates for (B) $\delta=4$ (“chaotic”), (C) $\delta=5.99$, and (D) $\delta=6.66$ (random walk), respectively.
@@
We repeated the previously depicted procedure for a sequence of values of the decay length in the range $[3-10]$ mm and measured the associated scaling exponents. The values of $\alpha(d)$ for different realizations are shown in Fig. 2 as blue points. The code to reproduce the simulation and the 3D model files (gltf) are available at the repository https://github.com/ggosti/HopBrain. This code also shows how to generate 3D maps for the Schaefer parcellation’s nodes activation states inspection, and how to export gltf files that can be viewed in immersive environments using open source webXR apps such as Aton which are part of the H2IOSC open cloud (Fanini et al., 2021; Fanini and Gosti, 2024) (Fig. 2B-D). Figure 2B-C can be accessed with a web-browser respectively with the following urls: https://aton.ispc.cnr.it/s/ggosti/20240612-6zkx4gyvi for $\delta=4$ (“chaotic”), https://aton.ispc.cnr.it/s/ggosti/20240612-ebn1rod5p for $\delta=5.99$, and https://aton.ispc.cnr.it/s/ggosti/20240612-a8e3initf for $\delta=6.66$ (random walk).